OpenStudy (anonymous):

The equation x2 + y2 − 2x + 2y − 1 = 0 is the general form of the equation of a circle. What is the standard form of the equation? (x − 1)2 − (y − 1)2 = 1

OpenStudy (anonymous):

the standard form of a circle would not have the minus sign between the two squares

OpenStudy (anonymous):

(x − 1)2 − (y − 1)2 = 1 (x − 1)2 + (y + 1)2 = 3 (x + 1)2 + (y − 1)2 = 1 (x − 1)2 + (y + 1)2 = 1

OpenStudy (anonymous):

those are the answer choices

OpenStudy (mathmale):

did you intend that (x − 1)2 − (y − 1)2 = 1 would represent one of the four answer possibilities?

OpenStudy (anonymous):

no, that was one of the answer choices

OpenStudy (mathmale):

You are likely to convey your math to others (and perhaps to yourself as well) if you'd please type the square of (x-1) as (x-1)^2 OR \[(x-1)^2\]OR

OpenStudy (mathmale):

OK: now, using Satellite73's obstervation, eliminate the answers that could not be correct. Then think about how to narrow down further to obtain the correct answer. strong hint: use "completing the square" to help y ourself identify the center (h,k) of this circle.

OpenStudy (anonymous):

it's either C or D

OpenStudy (mathmale):

As always I'm going to ask you to justify (give reasons for) your answer(s).

OpenStudy (anonymous):

may i please just have the answer